Database last updated on Sun, 01 Feb 2026 20:10:55 +0000 by localhost

Instrument constant changes for the week before 2024029 (4 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 11 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
085202402914619.1

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
0852024029126.6

Non-critical warnings from b-files' comments for days between 2024029 and 2024029 (1 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
ab_sk:HOME key pressed299B02924.085TORON#08510:27:081-1
dsp:Doing dispersion test with Hg (UV port)499B02924.085TORON#08510:48:141-1
checkhg:Last HG had to adjust the micrometer by at least 2399B02924.085TORON#08512:15:541-1
fr:Micrometer #1 reference found at step 9646 (+CAL)= 10655 Now set to o3 offset = 7832299B02924.085TORON#08512:59:121065510009000
ze:Zenith discrepancy = 76499B02924.085TORON#08520:13:1076-1010
pd:Going to menu199B02924.085TORON#08521:06:061-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssl
029TORONTO#085inside,menu,Must do HG now,Waiting until ZA = 170 now 130.5,Waiting until ZA = 170 now 132.2,Waiting until ZA = 170 now 134,Waiting until ZA = 170 now 135.6,Waiting until ZA = 170 now 137.3,Waiting until ZA = 170 now 138.9,Waiting until ZA = 170 now 140.5,Waiting until ZA = 170 now 142.1,Waiting until ZA = 170 now 143.6,Waiting until ZA = 170 now 145,Waiting until ZA = 170 now 146.4,Waiting until ZA = 170 now 147.7,Waiting until ZA = 170 now 148.9,Waiting until ZA = 170 now 150,Waiting until ZA = 170 now 151,Waiting until ZA = 170 now 151.9,Waiting until ZA = 170 now 152.6,Waiting until ZA = 170 now 153.2,Waiting until ZA = 170 now 153.6,Waiting until ZA = 170 now 153.9,Waiting until ZA = 170 now 15442644844
(1 sec)

Dead time (1 sec)

File nameCountSerialExpected (+/-5) Measured 30-day history
B02924.0854#0852830